OpenStudy (anonymous):

Which expression is always equivalent to v? v over five times five v over five times ten v over five times one v over ten times five

OpenStudy (anonymous):

Write an expression for each one and bear in mind that as long as x is not 0, then x/x = 1 and 1 is the multiplicative identity. That means that (m)(1) = m. Anything multiplied by 1 stays the same.

OpenStudy (anonymous):

So it's the last one

OpenStudy (anonymous):

Show some work and then I can say yes or no about that.

OpenStudy (anonymous):



OpenStudy (anonymous):

Does \[\frac{ v }{ 10 } \times 5 = \frac{ 5v }{ 10 }\] ?

OpenStudy (anonymous):

v over five times five v over five times ten v over five times one v over ten times five I need an answer quick.

OpenStudy (anonymous):

Why do you need an answer quick? Is this for a test?

OpenStudy (anonymous):

No I'm studying for finals and they gave me a sheet that helps me practice for the finals test. It's homework but I need the answer quick.

OpenStudy (anonymous):

It's v over 5 times 1

OpenStudy (anonymous):

We can't give just answers. We are not allowed, but even if we were allowed, a good tutor would never do that. You have to be actively involved or you will learn nothing. I repeat: you will learn nothing. And the whole goal of learning and this website is to get you to think and do your own work. I gave you some solid help and I will continue to do so, but you have to get involved. Another hint: factor top and bottom.